OpenStudy (anonymous):

Select the ordered pair from the choices below that is a solution to the following system of equations: 4y = x − 13 2x + 16y = 20 a. (14, 1/4) b. (-3, 1) c. (12, -1/4) d. (5, -2)

OpenStudy (cwrw238):

try replacing each value of x any into the equations a little help - its either a or c

OpenStudy (anonymous):

A ?

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

Just do trial and error - for each choice, plug in x and y into your system of equations and check if it gives you the right answer. For (A), plugging in x=14 and y=1/4 into the system of equations yields: \[4(1/4)=(14)-13\]\[1=1\] So that works for the first equation. For the second equation:\[2(14)+16(1/4)=20\]\[28+4=20\]\[32\neq20\] so it doesn't work for the second equation, thus (A) can't be the answer since it doesn't satisfy BOTH equations.